OpenStudy (anonymous):

Please help me? is my answer right? (or left? ;) 5^3√96

OpenStudy (anonymous):

Are you trying to simplify this equation?

OpenStudy (anonymous):

10 ^3 square root 12

OpenStudy (anonymous):

What are you trying to do? What's the question?

OpenStudy (anonymous):

To simplify

OpenStudy (anonymous):

My apologies for not saying sooner

hartnn (hartnn):

thats correct .

OpenStudy (anonymous):

you can simplify "square root 12" more...

OpenStudy (anonymous):

What's correct?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

10 ^3 square root 12 but as said, square root 12 can be simplified more

OpenStudy (anonymous):

how so?

hartnn (hartnn):

5^3√96=10 ^3 square root 12 is correct

hartnn (hartnn):

12=4*3

hartnn (hartnn):

\(\sqrt 12=\sqrt4\sqrt3=?\)

OpenStudy (anonymous):

I understand now; you're so kind, thank you.......................!

hartnn (hartnn):

ok, welcome ^_^